OpenStudy (anonymous):

What is the inverse of y = 2x2 – 5?

OpenStudy (hoblos):

2x2 = y+5 x2 = y/2 + 5/2 x = sqrt(y/2 + 5/2)

OpenStudy (anonymous):

if it is \[y=2x^2-5\] it does not have an inverse because it is not one to one. if you want to solve for x you can write \[y-5=2x^2\] \[\frac{y-5}{2}=x^2\] \[x=\pm\sqrt{\frac{y-5}{2}}\]